# A2VIS: Amodal-Aware Approach to Video Instance Segmentation

### v1

@@ Abstract
[…]
[inst1]organization=Department of Computer Science and Computer Engineering,addressline=1 University of Arkansas, city=Fayetteville, postcode=72701, state=AR, country=USA
[…]
[inst2]organization=Coupang, Inc,city=Seattle, postcode=98101, state=WA, country=USA
{graphicalabstract}
Figure 1: Overal architecture of the proposed A2VIS.
{highlights}
@@ 1 Introduction
Recent methods in VIS or MOT have offered various strategies to mitigate occlusion challenges such as learning inter-clip associations [12], leveraging motion features [25], or employing tracking queries [36]. While these methods signify progress, they primarily hinge on processing visible elements, often overlooking the comprehensive understanding of objects when parts are occluded. Higher frequencies of crossover and occlusion increase the likelihood of object identity switches. Therefore, refining the granularity of representation could be advantageous [27].
@@
Incorporating amodal representation into VIS is not straightforward, as we encounter two significant challenges: (i) effectively predicting amodal segmentation for each frame and (ii) maintaining consistent object tracking throughout the video. Regarding the first challenge, previous studies [35, 6] highlight the complexity of amodal segmentation, necessitating prior knowledge. Recent work attempts to model prior knowledge as shape prior [33, 7, 29]. However, these methods rely on pre-training with multiple shapes of specific object types, making them dependent on type priors and difficult to generalize. To resolve that, we explore the spatiotemporal prior knowledge (SaVos [35]), which build dense object motion across frames to explain amodal representation. To this end, we introduce a Spatiotemporal-prior Amodal Mask Head (SAMH) for amodal mask prediction. Intuitively, SAMH uses two types of spatiotemporal information: short-range and long-range. Short-range information is derived from visible segments in adjacent frames. If a portion of an object is obscured in one frame, it may become visible in a neighboring frame. Long-range information involves the amodal segmentation of the object across the entire video, which is useful when an object is heavily occluded for an extended sequence of frames. To achieve this, we model these two spatiotemporal priors using a masked attention mechanism, employing a visible spatiotemporal-prior mask (VSPM) for short-range information and an amodal spatiotemporal-prior mask (ASPM) for long-range information. These are further elaborated in the methods section.
@@ 2.1 Amodal Segmentation:
Amodal segmentation involves predicting an object’s shape, including both its visible and occluded parts, across both images and videos. While image-based amodal segmentation is usually straightforward by incorporating an occluded mask prediction ORCNN [8] or transformer-based mask head AISFormer [30], video-based amodal segmentation is more complex due to temporal consistency constraints. Approaches like SaVos [35] learn amodal representation by using visible parts from each frame and motion information via LSTM, while EoRaS [7] leverages the multi-layer view fusion and temporal information to address amodal video segmentation. Unlike existing video-based amodal segmentation approaches, which extract amodal video representation from visible masks in a multi-stage framework, A2VIS is an end-to-end framework that simultaneously detect, track, visible segmentation, and amodal segmentation for objects in videos.
@@ 2.2 Video Instance Segmentation (VIS):
Early VIS works like MaskTrack R-CNN [34], SIPMask [2], SGNet [20] extends image-based models Mask R-CNN [10] to videos by predicting frame-independent outputs and making association using post-processing during the inference stage. Later methods like IFC [16], Mask2Former-VIS [4], IDOL [32], SeqFormer [31], MinVIS [15], DVIS [37] take clip-level input and run sequentially with association algorithm during post-processing. Recently, VITA [13] introduces instance prototypes for video representation. Due to the emergence of long video benchmarks (OVIS [24]), those existing methods such as IFC, Mask2Former, or SeqFormer are limited in handling those benchmarks in an end-to-end manner. Lately, GenVIS [12] addresses this by extending VITA’s hypothesis with inter-clip association and criterion on instance prototypes. GenVIS utilizes a memory-based method, maintaining a single memory bank that accumulates all instance prototypes from processed clips. In contrast, A2VIS introduces a global-local instance prototype strategy. These prototypes are spatiotemporally decoded in each clip via SAMH module, enabling robust object associations, improving occlusion handling.
@@ 3.7 Classification Head
Table 1: VIS tracking comparison on FISHBOWL and SAILVOS using ResNet-50 and Swin-L backbones.
@@
Online MinVIS ResNet-50 37.12 25.13 41.76 48.90 3462 18.63 15.06 25.43 22.76 18452
DVIS ResNet-50 39.12 26.11 43.07 49.15 3811 20.34 15.28 28.02 23.76 17332
IDOL ResNet-50 39.93 26.53 42.91 49.14 3901 21.37 15.32 28.01 24.43 17232
IDOL Swin-L 41.22 28.47 48.74 55.23 3010 23.94 15.94 31.11 26.18 16660
Offline/Semi-onl. SeqFormer ResNet-50 36.81 25.23 41.09 48.62 3528 17.52 15.12 24.90 23.13 19121
Mask2Former-VIS ResNet-50 36.17 25.16 39.96 47.97 3952 17.44 14.92 25.55 22.12 19301
VITA ResNet-50 38.15 27.34 40.81 46.38 3820 18.32 15.09 26.54 23.58 18234
GenVIS ResNet-50 40.04 26.09 44.08 50.08 3480 21.89 15.41 27.93 24.78 18037
@@
Table 2: Amodal VIS tracking comparison on FISHBOWL and SAILVOS using ResNet-50 and Swin-L backbones.
@@ 4.3 Baselines
VIS Baselines. We compare A2VIS against SOTA VIS approaches to assess its performance in simultaneously detecting, segmenting, and tracking objects. We include both online methods such as IDOL, MinVIS [15], and DVIS, and offline/semi-online methods like SeqFormer, Mask2Former-VIS , VITA [13], and GenVIS [12], using both ResNet50 [11] and Swin-L [21] backbone networks, on FISHBOWL and SAIL-VOS datasets.
@@ 4.4 Quantitative Performance Comparison
Table 3: Tracking performance in comparison with MOT methods on FISHBOWL and SAILVOS using ResNet-50 backbone. All metrics are evaluated on amodal boxes.
@@ 4.8 Model complexity
To benchmark computational complexity, we conducted inference on an RTX 8000 GPU using the Swin-L backbone across 20 test videos of FISHBOWL. A2VIS, with 222.8M parameters, averaged 0.77 FPS, which is competitive with GenVIS-Amodal’s 220.3M parameters at 0.82 FPS. Despite the competitive complexity, A2VIS shows significant gaps in performance in comparison with GenVIS-Amodal.
Table 10: Model complexity comparison
@@
GenVIS-Amodal 40.66 3242 220.3M 0.82
A2VIS (Our) 43.08 2547 222.8M 0.77
4.8.1 Evaluation on real-world dataset

### v2

@@ 1 Introduction
Recent methods in VIS or MOT have offered various strategies to mitigate occlusion challenges such as learning inter-clip associations [14], leveraging motion features [32], [46],employing tracking queries [44], [46], or target-based unified approach [1]. While these methods signify progress, they primarily hinge on processing visible elements, often overlooking the comprehensive understanding of objects when parts are occluded. Higher frequencies of crossover and occlusion increase the likelihood of object identity switches. Therefore, refining the granularity of representation could be advantageous [34].
@@
SAILVOS [16] is the first paper to propose a dataset for amodal video instance segmentation. However, their proposed method on the dataset is limited to images, not videos. Specifically, their MaskJoint method is an extension of MaskRCNN [12], featuring two mask heads—one for visible mask prediction and another for amodal prediction. While this method introduces the concept of amodal segmentation, it does not fully integrate amodal segmentation into the video instance segmentation (VIS) problem. Incorporating amodal representation into VIS is not straightforward, as we encounter two significant challenges: (i) effectively predicting amodal segmentation for each frame and (ii) maintaining consistent object tracking throughout the video. Regarding the first challenge, previous studies [43, 8] highlight the complexity of amodal segmentation, necessitating prior knowledge. Recent work attempts to model prior knowledge as shape prior [41, 9, 36]. However, these methods rely on pre-training with multiple shapes of specific object types, making them dependent on type priors and difficult to generalize. To resolve that, we explore the spatiotemporal prior knowledge (SaVos [43]), which build dense object motion across frames to explain amodal representation. To this end, we introduce a Spatiotemporal-prior Amodal Mask Head (SAMH) for amodal mask prediction. Intuitively, SAMH uses two types of spatiotemporal information: short-range and long-range. Short-range information is derived from visible segments in adjacent frames. If a portion of an object is obscured in one frame, it may become visible in a neighboring frame. Long-range information involves the amodal segmentation of the object across the entire video, which is useful when an object is heavily occluded for an extended sequence of frames. To achieve this, we model these two spatiotemporal priors using a masked attention mechanism, employing a visible spatiotemporal-prior mask (VSPM) for short-range information and an amodal spatiotemporal-prior mask (ASPM) for long-range information. These are further elaborated in the methods section.
@@ 2.1 Amodal Segmentation:
Amodal segmentation involves predicting an object’s shape, including both its visible and occluded parts, across both images and videos. While image-based amodal segmentation is usually straightforward by incorporating an occluded mask prediction ORCNN [10], transformer-based mask head AISFormer [37], or amodal-box expansion [24], video-based amodal segmentation is more complex due to temporal consistency constraints. Approaches like SaVos [43] learn amodal representation by using visible parts from each frame and motion information via LSTM, while EoRaS [9] leverages the multi-layer view fusion and temporal information to address amodal video segmentation. Recent literature has seen the emergence of diffusion models for image-based amodal segmentation. Studies such as [29, 45] leverage pretrained diffusion models for inpainting tasks to enhance the completion of occluded regions.. Unlike existing video-based amodal segmentation approaches, which extract amodal video representation from visible masks in a multi-stage framework, A2VIS is an end-to-end framework that simultaneously detect, track, visible segmentation, and amodal segmentation for objects in videos.
@@ 2.2 Video Instance Segmentation (VIS):
Early VIS works like MaskTrack R-CNN [42], SIPMask [4], SGNet [23] extends image-based models Mask R-CNN [12] to videos by predicting frame-independent outputs and making association using post-processing during the inference stage. Later methods like IFC [18], Mask2Former-VIS [6], IDOL [40], SeqFormer [39], MinVIS [17], DVIS [46] take clip-level input and run sequentially with association algorithm during post-processing. Recently, VITA [15] introduces instance prototypes for video representation. Due to the emergence of long video benchmarks (OVIS [30]), those existing methods such as IFC, Mask2Former, or SeqFormer are limited in handling those benchmarks in an end-to-end manner. VideoCutler [28] introduces an unsupervised approach that achieves instance segmentation without relying on labeled data. OV-VIS [38] improve open-vocabulary VIS with higher speed but maintain accuarcy. [20] present offline-to-online knowledge dis- tillation (OOKD) for video instance segmentation (VIS), which transfers a wealth of video knowledge from an offline model to an online model for consistent prediction. TARVIS [1] presents a unified target-based segmentation approach, improves adaptability across different scenarios. Lately, GenVIS [14] addresses this by extending VITA’s hypothesis with inter-clip association and criterion on instance prototypes. GenVIS utilizes a memory-based method, maintaining a single memory bank that accumulates all instance prototypes from processed clips. In contrast, A2VIS introduces a global-local instance prototype strategy. These prototypes are spatiotemporally decoded in each clip via SAMH module, enabling robust object associations, improving occlusion handling.
@@ 3.7 Classification Head
Table 1: VIS tracking comparison on FISHBOWL and SAILVOS using ResNet-50 and Swin-L backbones. For each backbone, the best results are in bold, and the second-best results are underlined.
@@
Online MinVIS [17] ResNet-50 37.12 25.13 41.76 48.90 3462 18.63 15.06 25.43 22.76 18452
DVIS [46] ResNet-50 39.12 26.11 43.07 49.15 3811 20.34 15.28 28.02 23.76 17332
STEMSeg [2] ResNet-50 37.36 25.42 41.58 48.67 3624 18.89 15.28 25.67 22.95 18625
HEVIS [31] ResNet-50 37.47 25.33 41.42 48.85 3691 18.74 15.45 25.49 22.79 18814
TarVIS [1] ResNet-50 39.28 25.89 42.85 48.92 3956 20.12 15.47 27.79 23.95 17521
IDOL [40] ResNet-50 39.93 26.53 42.91 49.14 3901 21.37 15.32 28.01 24.43 17232
IDOL [40] Swin-L 41.22 28.47 48.74 55.23 3010 23.94 15.94 31.11 26.18 16660
Offline/Semi-onl. SeqFormer [39] ResNet-50 36.81 25.23 41.09 48.62 3528 17.52 15.12 24.90 23.13 19121
Mask2Former-VIS [6] ResNet-50 36.17 25.16 39.96 47.97 3952 17.44 14.92 25.55 22.12 19301
VITA [15] ResNet-50 38.15 27.34 40.81 46.38 3820 18.32 15.09 26.54 23.58 18234
GenVIS [14] ResNet-50 40.04 26.09 44.08 50.08 3480 21.89 15.41 27.93 24.78 18037
@@
Table 2: Amodal VIS tracking comparison on FISHBOWL and SAILVOS using ResNet-50 and Swin-L backbones. For each backbone, the best results are in bold, and the second-best results are underlined.
@@ 4.3 Baselines
VIS Baselines. We compare A2VIS against SOTA VIS approaches to assess its performance in simultaneously detecting, segmenting, and tracking objects. We include both online methods such as IDOL[40], MinVIS [17], StemSeg [2], TarVIS [1], HEVIS [31], and DVIS [46], and offline/semi-online methods like SeqFormer, Mask2Former-VIS , VITA [15], and GenVIS [14], using both ResNet50 [13] and Swin-L [25] backbone networks, on FISHBOWL and SAIL-VOS datasets.
@@ 4.4 Quantitative Performance Comparison
Table 3: Tracking performance in comparison with MOT methods on FISHBOWL and SAILVOS using ResNet-50 backbone. All metrics are evaluated on amodal boxes. Best results are in bold, and the second- best results are underlined.
@@ 4.6.4 Number of convolutional layers of Amodal Feature Extraction
4.6.5 Impact of SAMH on VIS benchmark
Table 8 presents the impact of SAMH on the VIS benchmark using the ResNet-50 model, comparing its performance on two datasets: FISHBOWL and SAILVOS. The results show that applying SAMH consistently improves performance across both datasets regarding VIS benchmark. For FISHBOWL, AP increases from 39.94 to 41.77, AR from 26.22 to 28.07, and IDF1 from 49.98 to 52.14, while the number of identity switches (IDs) decreases from 3493 to 3392, indicating better object tracking. Similarly, in SAILVOS, AP improves from 22.06 to 23.12, HOTA from 28.04 to 30.04, and IDF1 from 24.67 to 25.94, while the IDs decrease from 18142 to 17004, further demonstrating SAMH’s positive impact on tracking accuracy and stability. In general, these results highlight that SAMH improves both visible segmentation and tracking performance across different datasets.
Table 8: Impact of SAMH on VIS benchmark using ResNet-50
FISHBOWL SAILVOS
[…]
wo/ SAMH 39.94 26.22 43.91 49.98 3493 22.06 15.27 28.04 24.67 18142
w/ SAMH 41.77 28.07 46.12 52.14 3392 23.12 15.87 30.04 25.94 17004
4.6.6 Impact of different occlusion levels
Table 9 presents the performance of two methods, GenVis-Amodal and A2VIS, under two different occlusion rates: less than 50% and greater than 50%. For occlusion rates less than 50%, A2VIS outperforms GenVis-Amodal across all metrics, with an AP of 42.33 compared to GenVis-Amodal’s 38.62, and similarly higher values for AR, HOTA, and IDF1. A2VIS also has fewer identity switches (3123 vs. 3195 for GenVis-Amodal).
In scenarios with occlusion rates greater than 50%, A2VIS continues to show better performance than GenVis-Amodal, although the difference in scores is less pronounced. A2VIS achieves an AP of 33.14, while GenVis-Amodal reaches 29.78. Similar trends are observed for the other metrics, with A2VIS showing higher AR, HOTA, and IDF1 scores, and fewer identity switches. Overall, A2VIS consistently outperforms GenVis-Amodal, particularly under lower occlusion rates.
Table 9: Impact of different occlusion levels on FISHBOWL, using ResNet-50
[…]
$<$50% GenVis-Amodal 38.62 28.47 50.23 57.34 3195
A2VIS 42.33 29.24 52.18 60.12 3123
$>$50% GenVis-Amodal 29.78 22.65 39.87 49.92 3756
A2VIS 33.14 24.35 42.96 53.29 3657
4.7 Video Amodal Segmentation Comparison
@@ 4.8 Model complexity
To benchmark computational complexity, we conducted inference on an RTX 8000 GPU using the Swin-L backbone across 20 test videos of FISHBOWL. A2VIS, with 222.8M parameters, averaged 0.77 FPS, which is competitive with GenVIS-Amodal’s 220.3M parameters at 0.82 FPS. As in Table 12, despite the competitive complexity, A2VIS shows significant gaps in performance in comparison with GenVIS-Amodal.
Table 12: Model comparison between GenVIS-Amodal and our proposed A2VIS on tasks support, amodal performance and complexity.
Methods Tasks Performance Computational cost
@@
GenVIS-Amodal Amodal Segmentaiton 40.66 3242 220.3M 0.82
A2VIS (Our) Amodal Segmentaiton & Visible Segmentaiton 43.08 2547 222.8M 0.77
4.9 Evaluation on real-world dataset
@@ 5 Conclusion and Discussion
Limitation: A2VIS attempts to reconstruct the occluded regions using visible cues from adjacent frames. Thus, objects undergoing large intrinsic shape changes are less suitable for A2VIS. Moreover, our method focuses on handling in-frame occlusions only. In particular, our approach does not explicitly account for objects that are occluded by being partly or completely out of the frame or disappear in one frame and reappear in another. This limitation arises because existing amodal video instance segmentation datasets, such as FISHBOWL and SAILVOS, do not provide ground-truth annotations for objects that move out of the frame. As a result, we confine the amodal mask within the frame size.